it is given that f(x)=tan^-1 2x and g(x) =p tan^-1 x, where p is a contant. find the value of p for which f(1/2)=g(1/2)
1
Expert's answer
2020-06-15T19:18:39-0400

f(x) = tan-12x and g(x) = p tan-1x

Given that f("\frac {1}{2}" ) = g("\frac {1}{2}" )

=> tan-1(2*"\frac {1}{2}" ) = p tan-.1 ("\frac {1}{2}" )

=> tan-11 = p cot-12

So p cot-12 = tan-11

=> p cot-12 = "\frac {π}{4}"

=> p = "\frac {π}{4cot^{-1}2}"

So value of p is "\frac {π}{4cot^{-1}2}" which is approximately 1.69395